Topic: pH and Ka Value

Topic: pH and Ka Value Note: y-intercept =

Note:

y-intercept = 8.46

Henderson-Hasselbalch equation : pH = pKa + log[A-/HA]

Literature value of pKa = 9.245

Question: Use Henderson-Hasselbalch equation and y-intercept to determine the Ka value for the acid in the conjugate pair (NH3 and NH4Cl). Show your work.
Answer #1

Answer – We are given the Henderson-Hasselbalch equation and graph between the pH vrs log [NH3]/[NH4Cl]

We know the Henderson-Hasselbalch equation-

pH = pKa + log[A-/HA]

From the graph , y-intercept = 8.46

When we draw the graph between the pH vrs log salt/acid then there is straight line equation and the intercept is the pKa values.

So,    y-intercept = pKa = 8.46 when graph between the pH vrs log salt/acid.

We know,

pKa = - log Ka

so, Ka = 10-pKa

           = 10-8.46

           = 3.46*10-9

The actual Ka from pKa = 9.245 is 5.68*10-10, so there is maybe some error in the calculation or recording data.